A particle moves in a straight line with an initial

velocity of 30 m/s and constant acceleration 30 m/s2

.

(a) What is its displacement at t = 5 s? (b) What is its

velocity at this same time?


1
Expert's answer
2022-03-20T18:47:31-0400

(a) What is its displacement at t = 5 s?

"d=v_0t+at^2/2\\
=30\times 5+30\times 5^2/2=525\:\rm m"

(b) What is its velocity at this same time?

"v=v_0+at\\
=30+30\times 5=180\:\rm m/s"